how do i make this a function of x??
-2x+5y-6=-11

Answer by checkley77(12844)

-2x+5y-6=-11
-2X=-5Y-11+6
-2X=-5Y-5
X=-5Y/2-5/2
X=-2.5Y-2.5
X=-2.5(Y+1)